Prepositions in the four versions of Cursor Mundi

(How they support the typology of the MSS)

Judit Farkas

If the four edited versions of the Cursor Mundi are read parallelly, it is very much apparent that three versions are very similar linguistically, while the fourth one is much different. It is a well established supposition that the original Cursor MS was compiled somewhere in northern England. Why is this difference then?

In the Early English Text Society (EETS) edition of the four texts three studies can be read about the filiation of the MSS by Rev. Richard Morris (1891 [1962]) and Dr. H. Hupe (1874a, b [1962]). Hupe also gives two charts to illustrate his theory. As the two authors are contradictory, who is more right?

Also the possible northern origin of the original aroused my interest. If it is true, it must show also Scandinavian influence. So the question arose: which text shows the most Scandinavian influence, and if any of them does, whose theory does it support, Hupe’s or Morris’?

To try to find an answer a minor grammatical category—the prepositions—were examined in the four texts parallelly. 1000 lines were chosen randomly from different parts of the texts and the prepositions counted in each of them. After this the focus fell on those which had a definite Scandinavian influence, or might have been influenced. To compare them in the texts a chart was prepared to show not only the types and numbers of such prepositions, but also their percentages related to each other.

Obviously, this examination relies only on one particular grammatical means and therefore cannot be considered “absolute”. In spite of this, as the following essay will show, this analysis gives answers to some of the above questions, reveals contradictions in earlier studies as well as adds some new aspects of the genealogy and relationship of the four texts.

In the Prologue (ll. 1–270) of the Cursor Mundi the author makes clear his aim with this work:

10 Judit Farkas

“Efter haly kyre state -Dis ilk bok is es translate In to Inglis tong to rede For the loue of Inglis lede, Inglis lede of Ingland, For the commun at understand. (ll. 231–36)

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .

To laud and Inglis man i spell

-Dat understandes �at i tell… ” (ll. 249–50)

Here we should have a closer look at the statements in line 236: “For the commun at understand” and line 250: “-Dat understandes �at i tell… ”, namely what is the language the Cursor-author speaks about here which people understand? According to Morris’ Preface to the EETS edition in 1874–93, it is “the speech of the north-county folks” (p. vii, l. 3). However, if we read carefully the lines in the Prologue, we will not get any closer hints than what I have quoted above, about this tongue. It is obvious, though, from the painstaking work of philologists (like Morris himself, Hupe, Klauza (1888, 1889), etc.) that the original of the Cursor Mundi was compiled in the northern dialectal territory. Hupe (1874a, b [1962]) examined also the relationships of the surviving MS-copies to one another and to the supposed original and tried to set up a typology of these MSS, which is more or less acceptable. He studied and compared the phonology, noun–adjective–verb morphology, the personal pronouns, the conjunctions, as well as the rhymes of the individual MSS and set up his typology on the basis of them. Thus, he arrived at the following dialect-distribution or generation of the MSS:

MS–C: source: North, early 14th c.

copy: North Lancashire, 15th c. MS–F: source: more towards the South, 14th c.

copy: Western part of the ancient archdiocese of York, early 15th c. MS–G: source: little more towards the East than C: East of the ancient

diocese of Lichfeld and Coventry

copy: North-West of the ancient diocese of Lincolnshire, early 15th c. MS–T: source: a more Northern source, which is not likely to have been

purely Northumbrian

copy: South-Midland and more towards the West: South of the ancient diocese of Hereford, early 15th c.

Prepositions in Cursor Mundi 11

Reading his essay, we cannot deny the relevance of such a typology. However, if we have a closer look at the compared fields, we definitely miss some important territories, which may reveal new aspects as well as contribute, support, or even contradict his statements. I, myself, miss very much a rather important grammatical category from his examination, viz., the prepositions. As the Cursor Mundi —original and copies—is dated to the second half of the 14th century, first part of the 15th century, i.e., well within the Middle English period, it is obvious, that prepositions played an increasingly significant role in morphology as well as in syntax parallel with the morphological simplification of the language. Besides, prepositions — being grammatical words rather than independent, meaningful units—may reveal a lot more about the genealogy of the MSS. Therefore, I thought of taking the four edited versions (Cotton-Vespasian: C, Fairfax: F, Göttingen: G, Trinity: T) and picked out some 1000 lines at random and tried to compare the use of prepositions in them. My focus of interest, however, fell on those prepositions which may support the northern or non-northern origin of each MS, or to put it in a more exact way: whether the use of prepositions in the texts support Hupe’s typology?

I have already mentioned that my choice of lines was random, however, I could not neglect the fact that some of the MSS have big gaps in their texts, so I tried to choose parts which have more or less full parallels in each version. Here follows the list of the examined lines:

Prologue: ll. 1–270

Of the Fall of Lucifer: ll. 411–510 Isaac, Esau and Jacob: ll. 3409–3910 Aaron’s Rod, and the Death of Moses: ll. 6881–6920 The Choice of David’s Successor: ll. 8331–8434

In Appendix 1 there is a full list of prepositions with their frequencies in each text, while in Appendix 2 there is a list of those prepositions which have some northern character. I am giving here their frequencies as well as its proportion to the whole number of prepositions in each text as well as the non-northern means of expressing the same relationship.

Let us embark on the detailed analysis of Appendix 2 and see what conclusion can be drawn from all these data.

At the very beginning of the analysis we have to establish, however, that the type of prepositions falling in the scope of our interest should be defined more exactly. Therefore, we can set up, at least, four different categories. In Category 1 we can collect all those prepositions which are clearly loan-prepositions from the Old Norse language.^{1} In Category 2 we can select prepositions whose function can be paralleled and supported by Nordic influence. In Category 3 are those prepositions which preserved a typically northern form (as opposed to other dialects), while in Category 4 we can select those ones whose pattern—and here we are facing the problem of compounding—follows a Nordic pattern rather than a native one.

As Appendix 2 shows we can find altogether 11 prepositions which can be distributed in the formerly mentioned four categories. As Appendix 2 also shows, their number differs in each MS.^{2} Though independently of the MSS, we can map-up their categorization in the following way:

Category 1: til(le), in-til(l), vn-til(l) Category 2: with/wid/wi�, at, for to, for til(l) Category 3: fra/fro Category 4: out of/vt-of, apon/opon, for-wit

In MS–C we have altogether 150 occurrences of these 11 prepositions (as this is, in fact, the only MS, where all of them occur). The total number of prepositions is 661, which means 29.69% of the prepositions. (We may say, that it is one-third of the whole number of prepositional items.)

The analysis is started with Category 2, because the constructions there are reinforced by Nordic patterns. The infinitive-marker can be for to (22), at (10), for till (1), which mean 33.67% of all the infinitive constructions in the text considering to (65) as well. This is Category 2 as these constructions are reinforced by Nordic patterns. for to: ‘in order to; to’. If we take for example Swedish, the parallelism

can very well be demonstrated: PrE: we went to buy something, whereas in Swedish we have: vi gick for att köpa någonting, which is equivalent with the for to construction (cf. Holm & Nylund Lindgren 1977). In other modern Gmc languages we cannot find the same structure.^{3}

at: ‘to’. The OED says that it was lost in Southern and Western dialects. Hoad’s etymological dictionary (1987) states that originally this is an Old Teutonic preposition denoting motion and position ‘towards, further’, but this preposition was lost in the above-mentioned dialects and was replaced by to, just like in the Scandinavian dialects the to form was lost and the at form preserved. This statement very much supports the Scandinavian influence in the English usage, especially as the infinitive marker in the present-day Scandinavian languages is at/att.^{4} for till: ‘to’. It can be fairly straightforwardly parallelled with for to in

function and construction. However, it resembles a bit more the Nordic pattern as instead of the native to we find Nordic till.

wyt: ‘with’ (59). We should add this preposition also here, because its meaning is also due to Nordic influence as opposed to the original sense of this preposition in OE.^{5}

In Category 1 we find til(l) (19), in-til(l) (2) and vn-til(l) (1), which make 22 instances altogether, presenting 33.92%, 28.88% and 16.66% of the total number of constructions respectively as opposed to native to (37), into (5), unto (5).

In Category 3 there is only one preposition: fro/fra (20) as opposed to other dialectal and more common forms: from (0).

In Category 4 we find apon/opon (10), for-wit ‘before’ (3), and out(e) of (3). Obviously, they cannot be called either loan or northern dialectal forms. The only reason I felt they have to be taken into account here, is the pattern of compounding they follow: adverb+preposition as opposed to the English way: preposition+adverb (Cf. without, within, etc.)

We can follow the same way in the other three versions of the text. In MS–F the total number of prepositions is 659, the number of “northern” types is 137, which gives 20.78% of the whole (more than one-fifth). Their distribution into categories is as follows:

Category 2: to express infinitive (of purpose): for to (18), for til(l) (3), at (6) contrasted with to (75) which forms 26.47% of all the infinitive constructions. Wit(h) appears 57 times in the present-day meaning.

Category 1: til(l) (15) opposed to to (29), which forms 34.09%, intil(l) (3) opposed to into (1), i.e., 75%, vn-til(l) (1) opposed to unto (6) forming 14.28%.

Category 3: fra (20) and no from form at all.

Category 4: a-pon, vpon (13) (also apon-to (1)), out of (3).

In MS–G the total number of prepositional constructions is 676, the number of the examined types is 135, which makes 19.97% of the whole. (Actually, there does not seem to be too big difference between MSS F and G.)

Category 2: for to (24), at (1) which form 23.36% of all the infinitive constructions opposed to the number of to (82). Wid/wi� occurs 58 times here.

Category 1: til(l) (15) opposed to to (46), which makes 24.59%, intill (1) opposed to into (7), i.e., 14.28%.

Category 3: fra (21) and no from at all. Category 4: apon/vp-on (11), apon to (1), and vte-of (3).

In MS–T the total number of prepositions is 646, and that of the examined types is 92, which gives 14.27% of the whole. (Here we can already see a significant decrease comparing this version to C, but even if we compare it to F or G the difference is spectacular and obvious.)

Category 2: for to (13) as opposed to to (99), i.e., 11.60% of the whole infinitive constructions. And, of course, we have wi�/with with its 49 occurrences.

Category 1: til(le) (4) as opposed to to (54), which makes only 6.03%.

Category 3: fro (15) as opposed to from (4) and it forms 78.94%. Category 4: vp-on (5) and out of (6).

If we compare our table of data either in percentage or in actual numbers with what we stated initially about the dialect distribution or typology of the MSS, we can conclude the following:

(1) Considering the total number of prepositions which may have a Scandinavian or, at least, northern influence in any way, we can see really significant difference only between MSS C and T, 58 in number and almost 10% in proportion. (I think, however, this latter data is more reliable for the simple reason, that even the total number of prepositional constructions differs a bit among the MSS, due to different syntactical structures or solutions to convey the same or similar message.)

(2) If the categories are taken one by one and the items compared, we can arrive at just the same result. MS–C and MS–T differ spectacularly from each other. Here especially Category 2, the means of expressing infinitive, deserves special interest. In MS–C, the northern or Scandinavian type of infinitive particle gives one-third of the total constructions, which also means that it is more than half of the to-infinitive constructions. In MS–T, it is only ca. one-tenth of the whole and their numbers are considerably different. In MS–T, also the at and for-till forms are completely missing. And this is the only MS in which the from form appears as opposed to the fro/fra northern forms.

(3) What concerns MSS F and G, we cannot see really big differences between the two. We can take either the total proportion of the examined prepositions or the categories as well as the individual items, we get the same picture. Between the total numbers there is not entirely 1% (0.81%) difference, while between the infinitive markers not much more than 3% (3.11%).

(4) I do not want to go into the details of the other categories, because

they are interesting only in two respects:

(i) as they contribute their share to the total proportional analysis, and

(ii) their number/percentage in the individual MSS as opposed to more ‘native’ elements.

What conclusion can we draw from the above data?

(1) This analysis very well supports Morris’ and Hupe’s typology of the texts, that C, F, G are northern, while T is not.

(2) Where Morris’ and Hupe’s typologies contradict: Morris states, that text F has many peculiarities of the West-Midland dialect, while C and G are more characteristic specimens for the Northumbrian dialect. Hupe’s idea is that F must have been copied in the western part of the ancient diocese of York, which is clearly a more northern placement than Morris’. According to my analysis I must support Hupe here, as text F exhibits a little more northern characteristics than G.

(3) This analysis also seems to support Hupe’s supposition, that both F and G go back to a common source, from which they were copied, while C was copied from a completely different predecessor. What seems to me curious and — at least on the basis of this present study—not completely acceptable is that he originates T from the same source as G in both pedigrees he gives on pages 103 and 116 of his previously mentioned study, even though the latter one is a revised version of the former one. The only difference is between the two that in the first version (p. 103) he directly generates T from G, while in the latter figure (p. 116), he traces them back to the same source only. Of course, he examined much more aspects of the texts, than we here. Still, if we read the texts parallel line by line, it is very spectacular how different T usually is from the other three in many respects (spelling, sentence structures, morphological markers, choice of words, etc.). The only thing which seems to be true is that from among the 3 versions it is really text G it is closest to. Even though we have established and accepted the fact that text T was copied in a much more southern-ward area, and thus it reflects South-Midland characteristics, if it had gone back to the same source, especially if it had been copied from text G, it should not be so different from the other text. Therefore, this point needs to be more scrutinized.

appendix 1: Occurrence of Prepositions in the MSS

Preposition  MS–C MS–F MS–G MS–T
abote/abute/a-boute  5  6  5  5
agayn/a-gain/a-gains/a�eyn  7  7  8  9
amang/amange/emang/amone  1  1  1  1
apon to  1  1  1  0
apon/opon/vpon  10  13  11  5
at (adv.)  7  13  8  12
at (inf.)  10  6  1  0
be/bi/by  13  14  10  14
be-side/biside/bisyde  2  2  2  2
before/bifor(e)  0  3  2  1
bytwix/bitwix/-tuix/-twene/-twixe 6  5  6  3
bot/but  4  3  4  4
bot/but of  1  1  1  1
efter/ofter/aftir  6  6  6  6
for til(le)  1  3  0  0
for  33  32  29  26
for to  22  18  24  13
fro/fra  20  20  21  15
from  0  0  0  4
�onder  1  1  1  1
in til(l)  2  3  1  0
in to  5  1  7  7
in (time)  6  5  4  0
in (place)  73  74  77  73
of (adv.)  50  63  67  62
of (possessive)  140  129  129  126
on (time, place, manner)  21  15  16  22
on to/vnto  5  6  4  0
ouer  3  2  2  2
out(e) of/vte of  3  3  3  6

Preposition  MS–C MS–F MS–G MS–T
thoro/thoru/�orow/�our�e  4  5  3  3
til(l)  19  15  15  4
to (inf.)  65  75  82  99
to (adv.)  37  29  46  54
tofore  0  0  0  1
under/vndur/vnder/undir  2  2  2  1
vn-til(l)  1  1  0  0
vp to  0  0  1  0
wit-in/wi� in/wid-inne  2  2  1  1
witouten/wi�-/wid-  14  14  16  13
wyt/with/wit/wid/wi�  59  57  58  49
Total number of preps.: 661  659  676  646

appendix 2: Prepositions in Cursor Mundi (1000 lines)

MS–C MS–F MS–G MS–T
Total:  661  659  676  646
Scandinavian  150=  137=  135=  92=
influence:  29.69% 20.78% 19.97% 14.27%

MS–C:

Cat. 1:  til(l): 19=33.92% opposed to to: 37
in-till: 2=28.88% opposed to into: 5
vn-till: 1=16.66% opposed to vnto: 5
Cat. 2:  for-to: 22
at: 10
for-till: 1
Total: 33=33.67% opposed to to: 65
wyt: 59
Cat. 3:  fro/fra: 20  opposed to from: 0
Cat. 4: apon/opon: 10
for-wit: 3
out(e) of : 3

18 Judit Farkas

MS–F:

Cat. 1:  til(l): 15=34.09% opposed to to: 29
in-til(l): 3=75.00% opposed to into: 1
vn-til(l): 1=14.28% opposed to vnto: 6
Cat. 2:  for-to: 18
for-till: 3
at: 6
Total: 27=26.47% opposed to to: 75
wit(h): 57
Cat. 3: fra: 20 opposed to from: 0
Cat. 4: a-pon/vpon: 13
a-pon to: 1
out of : 3

MS–G:

Cat. 1:  til(l): 15=24.59% opposed to to: 46
intill: 1=14.28% opposed to into: 7
Cat. 2:  for to: 24
at: 1
Total: 25=23.36% opposed to to: 82
wid/wi�: 58
Cat. 3: fra: 21 opposed to from: 0
Cat. 4: apon/vp-on: 11
apon to: 1
vte-of : 3

MS–T:

Cat. 1:  til(le): 4= 6.03% opposed to to: 54
Cat. 2:  for to: 13=11.60% opposed to to: 99
wi�/with: 49
Cat. 3:  fro: 15=78.94% opposed to from: 4
Cat. 4:  vp-on: 5
out of : 6

notes

[1] I am using ‘Old Norse’ in the classical sense of the term, i.e., as it is used in more recent European philology as opposed to Old Swedish (cf. Wessén 1965, Gordon 1986). For the Old Norse data, I have been using Gordon 1986.

[2] This difference in number is partly due to the different wording of the MSS, but the question of meter cannot be neglected either. However, this latter is out of the scope of the present study. It is mentioned only to complete the picture.

[3] In present-day German we find a similar construction: um zu, but it is a much later development and does not follow the same word-pattern, only the function is similar.

[4] It is true, that the OED also says, that this pattern, and especially the ME usage, must have been reinforced by a similar French construction, and it was restricted to purpose-clauses, which is true for the previous pattern, too.

[5] The original meaning of this preposition is ‘against, in exchange for’, etc. The original preposition for expressing the present-day ‘togetherness, accompaniment’, etc. was mid. Actually mid was in force up till the 14th century by which time with pushed it out from this sense. One of the possible causes could be the similarity between English wi� and ON vi�, which had all the meanings of pre-English with, as well as the similarity between OE mid and Scandinavian med ‘with’.

references

Text edition: Morris, Richard, ed. (1962) Cursor Mundi. Early English Text Society 99. Oxford: OUP.

Gordon, E. V. (1986) An Introduction to Old Norse. Oxford: OUP.

Hoad, Terry F. (1987) The Concise Oxford Dictionary of English Etymology. Oxford: Clarendon Press.

Holm, Britta and Elizabeth Nylund Lindgren (1977) Deskriptive svensk grammatik. Trelleborg: Skriptor.

Hupe, H. (1874a [1962]) ‘On the filiation of the text of the MSS of the ME poem Cursor Mundi’. In Morris 1962 : 59–103.

Hupe, H. (1874b [1962]) ‘Cursor studies and criticism of the dialects of its MSS’. In Morris 1962 : 109–200.

Klauza, Max (1888) ‘Zum Handschriftverhältniss und zur Textkritik des Cursor Mundi [On the filiation of the MSS and on the textual criticism of Cursor Mundi]’. Englische Studien XI. 235–275.

20 Judit Farkas

Klauza, Max (1889) ‘Zu den Quellen und dem Handschriftverhältniss des Cursor Mundi [On the sources and the filiation of the MSS of Cursor Mundi]’. Englische Studien XII. 451–458.

Morris, Richard (1891 [1962]) ‘Preface’ to Cursor Mundi. In Morris 1962 : vii–xxx. OED = The Oxford English Dictionary (1953) Oxford: OUP. Wessén, Elias (1965) Svensk Språkhistoria III. Stockholm: Almquist & Wiksell.